MICROWAVE SPECTRUM AND GEOMETRY OF H_{3}P⋯AgI

\underline{N. R. WALKER}, S. L. STEPHENS, W. MIZUKAMI, D. P. TEW AND A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K..

The pure rotational spectrum of the vibrational ground state of H_{3}P⋯AgI has been measured by chirped-pulse FTMW spectroscopy. The complex is generated via laser ablation (532 nm) of a silver rod in the presence of CF_{3}I, PH_{3} and argon. It is subsequently stabilized and interrogated in the cold environment of a supersonic jet. The rotational constant, B_{0}, and the centrifugal distortion constant, D_{J} , have been measured for H_{3}P⋯^{107}AgI and H_{3}P⋯^{109}AgI. The spectrum of the complex is consistent with a C_{3v} geometry and a linear arrangement of the P, Ag and C atoms. The measured rotational constants allow a preliminary determination of the geometry of the molecule. The nuclear quadrupole coupling constant of the iodine atom, χ_{aa}(I) , is also established. The experimental results are compared with theory performed at the explicitly-correlated coupled-cluster singles, doubles and perturbative triples level.